# SnapGen-V: Generating a Five-Second Video within Five Seconds on a Mobile Device

### v1

@@ Abstract
We have witnessed the unprecedented success of diffusion-based video generation over the past year. Recently proposed models from the community have wielded the power to generate cinematic and high-resolution videos with smooth motions from arbitrary input prompts. However, as a supertask of image generation, video generation models require more computation and are thus hosted mostly on cloud servers, limiting broader adoption among content creators. In this work, we propose a comprehensive acceleration framework to bring the power of the large-scale video diffusion model to the hands of edge users. From the network architecture scope, we initialize from a compact image backbone and search out the design and arrangement of temporal layers to maximize hardware efficiency. In addition, we propose a dedicated adversarial fine-tuning algorithm for our efficient model and reduce the denoising steps to $4$. Our model, with only $0.6$B parameters, can generate a $5$-second video on an iPhone 16 PM within $5$ seconds. Compared to server-side models that take minutes on powerful GPUs to generate a single video, we accelerate the generation by magnitudes while delivering on-par quality.
@@ 2 Related Work
Video Diffusion Models. Denoising Diffusion Probabilistic Models [17] is the trending paradigm for building video diffusion models, demonstrating photorealistic quality and generic generation capabilities. Pioneer works often start from a pre-trained text-to-image diffusion model [45] and insert temporal layers to model motions along frame sequence [4, 62, 13, 3, 74]. In addition, training-free noise tuning techniques are proposed to ease the alignment between frames [63, 43, 23, 34]. Later, with the emergence of large-scale, high-quality video datasets [7, 38] and Transformer backbones [40], subsequent works curate their own dataset and build large video diffusion models with exceptional quality, such as the open-sourced CogVideoX [67], Mochi 1 [55], PyramidalFlow [21], Allegro [77], and close-sourced ones including Hailuo [37], Runway Gen 3 Alpha [46], Kling [24], Luma Dream Machine [1], Pika 1.5 [2], Sora [39], and MovieGen [41]. Remarkably, the open-sourced projects Open-Sora [76] and Open-Sora-Plan [26] provide the community with reliable implementations to replicate large-scale video diffusion models.
Dividing by task type, video diffusion models can be categorized into text-to-video generation [13, 18, 67, 55, 27, 28, 21, 77, 56, 5, 6, 42, 15], image-to-video generation [3, 78], or specific motion controls [71, 14, 44, 75, 61]. Though some work [29, 70, 8] aim to improve the efficiency of diffusion models, the acceleration for mobile deployment of video diffusion models is still in absent. Popular video models [67, 21, 55] can only run on server-level GPUs to generate videos in tens of seconds or even minutes.
@@ 3 Method
Our objective is to achieve high-fidelity and temporally consistent video generation on mobile devices. However, current text-to-video diffusion models face two key challenges in reaching this goal: (a) the memory and computation requirement is beyond the capability of even the most powerful mobile chips, i.e. iPhone A18 Pro, and (b) denoising with dozens of steps to generate a single output further slows down the process. To address these challenges, we propose a three-stage framework to accelerate video diffusion models on the mobile platform. First, we prune from a pre-trained text-to-image diffusion modelwto obtain an efficient spatial backbone. Second, we inflate the spatial backbone with a novel combination of temporal modules which are searched out with our mobile-oriented metrics. eFinally, through adversarial training, our efficient model attains the capability to generate high-quality videos in only $4$ steps.
@@ 3.1 Preliminaries
Following  [76], we employ a spatial-temporal VAE to compress image and video data into the latent space. Given video or image data $\mathbf{v}\in\mathbb{R}^{n\times 3\times H\times W}$, where $n$ is the number of frames with height $H$ and width $W$, the spatial-temporal encoder, $\mathbf{E}$, maps the data to a latent space. The encoded frames are represented as $\mathbf{x}_{0}=\mathbf{E}(\mathbf{v})$, resulting in $\mathbf{x}_{0}\in\mathbb{R}^{\tilde{n}\times 4\times\tilde{H}\times\tilde{W}}$. Here, $\mathbf{x}_{0}\sim p_{data}(\mathbf{x}_{0})$ is a $4$-channel latent. with a temporal compression of $\tilde{n}=n/4$, and spatial compression as $\tilde{H}=H/8$, $\tilde{W}=W/8$.
@@ 3.2 Hardware Efficient Model Design
Latency and Memory Guided Architecture Search. Prior to searching the architecture, we construct a look-up table containing the inference latency and the memory footprint of different temporal layers. For each candidate operator (i.e., ConvND, SelfAttentionND, CrossAttentionND), we benchmark the latency and memory consumption under different spatial-temporal resolutions on hardware. We then clean the search space by eliminating OOM states. Then we perform evolutionary search to obtain the temporal design with Pareto optimality. The architecture candidate is trained on precomputed video latents for $20K$ iterations with the spatial backbone frozen, and is evaluated on VBench [19] to obtain the scores as the quality metric. We include the detailed search algorithm, action space, and total search time in the supplementary material.
@@ 4.3 Ablation Analysis
w/ I Scaling $\Delta$AQ $\Delta$IQ $\Delta$DD $\Delta$MS $\Delta$Total
@@
Temporal Scaling ×$\times 0.5$ $-1.90$ $+0.04$ $-0.56$ $-0.61$ $-1.58$
×$\times 2$ $+0.37$ $+0.10$ $+0.56$ $+0.12$ $+0.27$
×$\times 4$ $+0.49$ $-0.17$ $+0.91$ $+0.17$ $+0.31$
Table 3: Analysis of training data scheme and quality of efficient architecture. The baseline model adopts the best suitable architecture and is trained with joint image-video datasets. The “scaling” indicates the number of the temporal layers comparing to the baseline. The “AQ”, “IQ”, “DD”, and “MS”, are aesthetic quality, image quality, dynamic degree, and motion smoothness in VBench [19]. The benchmark metrics are presented as the differences from the baseline model, where negative values indicates a decrease in performance over the baseline vice versa.
@@
Head w/ I $m$ TF MO Q S Total
@@
Table 5: Analysis of the number of inference steps. We measure VBench [19] score with different numbers of inference steps. In the results, “DD”, “OC”, and “AQ” denote the dynamic degree, object class, and aesthetic quality scores, respectively.
@@ Appendix D Qualitative Comparison Results
The comparison of our model with OpenSora-v1.3[76] and CogVideoX-2B [67] is shown in Fig. A2. More comparisons are presented in the first section of our accompanying webpage.
Appendix E More Qualitative Results
@@
To facilitate a more interactive and illustrative experience, these qualitative results are provided in video format. Readers are recommended to check these results in the second section of our accompanying webpage. This visualization provides a more nuanced understanding of the temporal and visual qualities of our method.
Appendix F Demo Settings
@@
Our demo is evaluated on an iPhone 16 Pro Max, equipped with an Apple A18 Pro chipset featuring a six-core CPU, six-core GPU, and 16-core Neural Engine. Our model is converted to FP16 and executed on the Neural Engine and the CPU cores. To enhance efficiency, timestep embeddings are also pre-computed since these values are fixed for each timestep. The inference pipeline takes four denoising steps without classifier-free guidance. To enable a fast mobile demo with pleasing quality, we adjust the input size for denoiser to $15\times 64\times 64$, which yields an output video clip with $51$ frames $512\times 512$ resolution. To ensure the video quality, the model is further finetuned with video datasets with a framerate of $10$ fps. Hence, the $51$ frame clip is $5.1$ seconds in length. Our UNet model is exported by CoreML and benchmarked using Xcode Performance tools. Furthermore, the exported model is split into two parts for loading and execution efficiency. The latency benchmark screenshots are shown in Fig. A5, thus one denoising step takes 1.02 seconds. Similarly, the text-encoder and VAE-decoder take 6 ms and 0.5 seconds, respectively. Thus, the entire inference pipeline takes less than 5 seconds on average. We exhibit the mobile demo screenshots in Fig. A4.

### v2

@@ Abstract
We have witnessed the unprecedented success of diffusion-based video generation over the past year. Recently proposed models from the community have wielded the power to generate cinematic and high-resolution videos with smooth motions from arbitrary input prompts. However, as a supertask of image generation, video generation models require more computation and are thus hosted mostly on cloud servers, limiting broader adoption among content creators. In this work, we propose a comprehensive acceleration framework to bring the power of the large-scale video diffusion model to the hands of edge users. From the network architecture scope, we initialize from a compact image backbone and search out the design and arrangement of temporal layers to maximize hardware efficiency. In addition, we propose a dedicated adversarial fine-tuning algorithm for our efficient model and reduce the denoising steps to $4$. Our model, with only $0.6$B parameters, can generate a $5$-second video on an iPhone 16 PM within $5$ seconds. Compared to server-side models that take minutes on powerful GPUs to generate a single video, we accelerate the generation by magnitudes while delivering on-par quality. Project page at https://snap-research.github.io/snapgen-v/.
@@ 2 Related Work
Video Diffusion Models. Denoising Diffusion Probabilistic Models [17] is the trending paradigm for building video diffusion models, demonstrating photorealistic quality and generic generation capabilities. Pioneer works often start from a pre-trained text-to-image diffusion model [45] and insert temporal layers to model motions along frame sequence [4, 62, 13, 3, 75]. In addition, training-free noise tuning techniques are proposed to ease the alignment between frames [63, 43, 23, 34, 73]. Later, with the emergence of large-scale, high-quality video datasets [7, 38] and Transformer backbones [40], subsequent works curate their own dataset and build large video diffusion models with exceptional quality, such as the open-sourced CogVideoX [67], Mochi 1 [55], PyramidalFlow [21], Allegro [78], and close-sourced ones including Hailuo [37], Runway Gen 3 Alpha [46], Kling [24], Luma Dream Machine [1], Pika 1.5 [2], Sora [39], and MovieGen [41]. Remarkably, the open-sourced projects Open-Sora [77] and Open-Sora-Plan [26] provide the community with reliable implementations to replicate large-scale video diffusion models.
Dividing by task type, video diffusion models can be categorized into text-to-video generation [13, 18, 67, 55, 27, 28, 21, 78, 56, 5, 6, 42, 15], image-to-video generation [3, 79], or specific motion controls [71, 14, 44, 76, 61]. Though some work [29, 70, 8, 73] aim to improve the efficiency of diffusion models, the acceleration for mobile deployment of video diffusion models is still in absent. Popular video models [67, 21, 55] can only run on server-level GPUs to generate videos in tens of seconds or even minutes.
@@ 3 Method
Our objective is to achieve high-fidelity and temporally consistent video generation on mobile devices. However, current text-to-video diffusion models face two key challenges in reaching this goal: (a) the memory and computation requirement is beyond the capability of even the most powerful mobile chips, i.e. iPhone A18 Pro, and (b) denoising with dozens of steps to generate a single output further slows down the process. To address these challenges, we propose a three-stage framework to accelerate video diffusion models on the mobile platform. First, we prune from a pre-trained text-to-image diffusion model to obtain an efficient spatial backbone. Second, we inflate the spatial backbone with a novel combination of temporal modules which are searched out with our mobile-oriented metrics. Finally, through adversarial training, our efficient model attains the capability to generate high-quality videos in only $4$ steps.
@@ 3.1 Preliminaries
Following  [77], we employ a spatial-temporal VAE to compress image and video data into the latent space. Given video or image data $\mathbf{v}\in\mathbb{R}^{n\times 3\times H\times W}$, where $n$ is the number of frames with height $H$ and width $W$, the spatial-temporal encoder, $\mathbf{E}$, maps the data to a latent space. The encoded frames are represented as $\mathbf{x}_{0}=\mathbf{E}(\mathbf{v})$, resulting in $\mathbf{x}_{0}\in\mathbb{R}^{\tilde{n}\times 4\times\tilde{H}\times\tilde{W}}$. Here, $\mathbf{x}_{0}\sim p_{data}(\mathbf{x}_{0})$ is a $4$-channel latent, with a temporal compression of $\tilde{n}=n/4$, and spatial compression as $\tilde{H}=H/8$, $\tilde{W}=W/8$.
@@ 3.2 Hardware Efficient Model Design
Latency and Memory Guided Architecture Search. Prior to searching the architecture, we construct a look-up table containing the inference latency and the memory footprint of different temporal layers. For each candidate 1D or 3D operator (i.e., ConvND, SelfAttentionND, CrossAttentionND), we benchmark the latency and memory consumption under different spatial-temporal resolutions on hardware. We then clean the search space by eliminating OOM states. Then we perform evolutionary search to obtain the temporal design with Pareto optimality. The architecture candidate is trained on precomputed video latents for $20K$ iterations with the spatial backbone frozen, and is evaluated on VBench [19] to obtain the scores as the quality metric. We include the detailed search algorithm, action space, and total search time in the supplementary material.
@@ 4.2 Quantitative Comparisons
User Study.  Human evaluations are conducted between our model and baselines as in the Tab. 3. We generate videos from VBench and Movie Gen Bench prompts and ask human labelers to pick the best results across prompt alignment, aesthetics, and motion. The result shows that our model outperforms baseline metrics by a large margin.
Win Rate Prompt Alignment Aesthetics Motion
OpenSora-1.2 24.9% 21.2% 20.4%
CogVideoX-2B 30.7% 30.5% 30.3%
Ours 44.4% 48.3% 49.3%
Table 3: User Study between OpenSora-1.2, CogVideoX-2B, and our model on prompt alignment, aesthetics, and motion.
@@ 4.3 Ablation Analysis
w/ I Scaling $\Delta$T $\Delta$AQ $\Delta$IQ $\Delta$DD $\Delta$MS $\Delta$Total
@@
Temporal Scaling ×$\times 0.5$ $-22\%$ $-1.90$ $+0.04$ $-0.56$ $-0.61$ $-1.58$
×$\times 2$ $+32\%$ $+0.37$ $+0.10$ $+0.56$ $+0.12$ $+0.27$
×$\times 4$ $+107\%$ $+0.49$ $-0.17$ $+0.91$ $+0.17$ $+0.31$
Table 4: Analysis of training data scheme, latency, and quality of efficient architecture. The baseline model adopts the best suitable architecture and is trained with joint image-video datasets. The “scaling", and “T" indicates the number of the temporal layers, and the latency comparing to the baseline. The “AQ", “IQ", “DD", and “MS", are aesthetic quality, image quality, dynamic degree, and motion smoothness in VBench [19]. The benchmark metrics are presented as the differences from the baseline model, where negative values indicates a decrease in performance over the baseline vice versa.
@@
Head w/ I w/ $\mathcal{L}_{\text{adv}}$ $m$ TF MO Q S Total
@@
Table 6: Analysis of the number of inference steps. We measure VBench [19] score with different numbers of inference steps. In the results, “DD", “OC", and “AQ" denote the dynamic degree, object class, and aesthetic quality scores, respectively.
@@ 5 Discussion and Conclusion
  • Effect of Adversarial Fine-tuning: We further discuss the effect of adversarial finetuning for step distillation in Appendix G.
  • Latency Analysis: Appendix H shows the latency analysis of different temporal blocks.
@@ Appendix A Search Algorithm
Figure A1: Comparison between the SD1.5 and our efficient spatial UNet backbone.
@@ Appendix C Qualitative and Quantitative Results for the Spatial Backbone
Additionally, we exhibit the quality comparison of our spatial backbone with SD1.5 in Fig. A1.
@@
Appendix D Qualitative Comparison Results
The comparison of our model with OpenSora-v1.3[77] and CogVideoX-2B [67] is shown in Fig. A4. More comparisons are presented in project page.
Appendix E More Qualitative Results
To facilitate a more interactive and illustrative experience, these qualitative results are provided in video format. Readers are recommended to check these results in project page. This visualization provides a more nuanced understanding of the temporal and visual qualities of our method.
Appendix F Demo Settings
Our demo is evaluated on an iPhone 16 Pro Max, equipped with an Apple A18 Pro chipset featuring a six-core CPU, six-core GPU, and 16-core Neural Engine. Our model is converted to FP16 and executed on the Neural Engine and the CPU cores. To enhance efficiency, timestep embeddings are also pre-computed since these values are fixed for each timestep. The inference pipeline takes four denoising steps without classifier-free guidance. To enable a fast mobile demo with pleasing quality, we adjust the input size for denoiser to $15\times 64\times 64$, which yields an output video clip with $51$ frames $512\times 512$ resolution. To ensure the video quality, the model is further finetuned with video datasets with a framerate of $10$ fps. Hence, the $51$ frame clip is $5.1$ seconds in length. Our UNet model is exported by CoreML and benchmarked using Xcode Performance tools. Furthermore, the exported model is split into two parts for loading and execution efficiency. The latency benchmark screenshots are shown in Fig. A7, thus one denoising step takes 1.02 seconds. Similarly, the text-encoder and VAE-decoder take 6 ms and 0.5 seconds, respectively. Thus, the entire inference pipeline takes less than 5 seconds on average. We exhibit the mobile demo screenshots in Fig. A6 and project page.
Appendix G Effect of Adversarial Fine-tuning
Tab. 5 also shows the effect of adversarial fine-tuning. Tuning without adversarial loss can not yield promising results compared to the baseline for step distillation.
Appendix H Latency Analysis
Figure A2: Latency
The latency of different temporal blocks is shown in Fig. A2.